# Optimizing entanglement in two-qubit systems

### v1

@@ Abstract
We investigate entanglement in two-qubit systems using a geometric representation based on the minimum of essential parameters. The states are X–shaped and host pairs of identical populations while reducing the number of coherences involved. A geometric $L$-measure of entanglement is introduced as the distance between the points that represent entangled states and the closest point that defines separable states. Our results give a geometric meaning to the Hill-Wootters concurrence $C$ since $L$ coincides with $C$. Furthermore, unlike $C$, the measure $L$ distinguishes the rank of states with maximum entanglement. We construct a time-dependent two-qubit state, traced out over the complementary components of a pure tetra-partite system, and find that its one-qubit states share the same entropy. The entanglement measure results bounded from above by the envelope of the minima of such entropy.
@@ 1 Main
The entanglement in a system can be characterized and quantified from different points of view [7, 8, 9, 10, 11, 12, 13]. However, the separability criterion introduced by Peres [14] and Horodecki’s [15] provides a powerful tool for its detection. In this work, the Peres-Horodecki (also known as positive partial transpose –PPT–) criterion is applied to construct a convex representation of two-qubit states that permits the study of entanglement in purely geometric form. It is well known that this criterion is necessary and sufficient to detect entanglement in $2\otimes 2$ and $2\otimes 3$ quantum systems [15], so we will use it to identify entanglement in two-qubit mixed states. There are several incentives to consider these states, mainly because generating pure entangled states is not an easy task, so entangled states could be prepared as statistical mixtures for purely experimental reasons, but in general they can be the result of a partial trace of the quantum state of a larger system.
@@
In general, mixed states characterize the interaction between the system and its environment, and the study of their entanglement properties is not only more complicated than that of pure states, but lesser understood. Although the first entangled system found in physics was in a pure state [2], it is in one of the Bell states [4], shortly after mixed states were identified that also disrupt the foundations of quantum theory [16]. These states, named after Werner, can be modeled using hidden variables without violating the Bell inequalities, and find applications in a diversity of quantum entanglement subjects [17, 18, 19, 20, 21, 22, 23]. Our results include a wide class of Werner states as a particular case.
@@
We investigate the entanglement properties of two-qubit systems by using the minimum of essential parameters. The idea is to construct density operators whose reduced one-qubit states share the same entropy, regardless of whether the state of the entire system is pure or mixed. In this way, we identify X–shaped states [24, 25, 26], denoted $\rho_{X}$, that have the capacity to host pairs of identical populations while reducing the number of coherences involved. Demanding positive semidefinite, self-adjoint and normalized density operators $\rho_{X}$, and imposing the Peres-Horodecki criterion, we find a system of inequalities that related matrix elements must satisfy to determine both, the eligibility of $\rho_{X}$ as a quantum state, and the possibility of finding entanglement. The solution of such a system defines a convex set $\mathcal{S}$ that can be divided in subsets whose points define separable states and regions where entanglement may be found.
@@
Based on the convex representation provided by $\mathcal{S}$, we introduce a geometric $L$-measure of entanglement, a distance-like entanglement measure [27], that coincides with the Hill-Wootters generalized concurrence $C$ [28, 29] when applied to states $\rho_{X}$. Thus, our results give $C$ a geometric meaning that is not obvious without the convex structure $\mathcal{S}$. Furthermore, unlike $C$, the measure $L$ distinguishes the rank of states with maximum entanglement. The latter is relevant to determine not only the amount of quantum correlations but rather the quantumness of the state [30], so it shows how much useful a state is for a specific quantum processing task [31, 32]. Another advantage of the geometric $L$-measure is that, with respect to other entanglement measures, it allows comparatively easy calculation.
@@
We also study what happens to entanglement when the system depends explicitly on time. Modeling an entangled diatomic system whose one-qubit parts are placed in independent, isolated and identical electromagnetic cavities, each of which contains $n$ photons, we construct a tetra-partite pure state that continues to evolve unitarily as a pure state. However, the reduced diatomic state does not evolve in a reversible way (systems like this may show signatures of chaos [33]), which makes it an excellent subject of study to understand the behavior of entanglement in systems described by time-dependent statistical mixtures. Regardless of the dynamical law obeyed by the time-dependent diatomic state, the corresponding one-qubit states share the same von Neumann entropy $S$, which oscillates as time passes. Our results verify that finding reduced states of maximum mixing does not provide information about the entanglement (if any) of the entire system when it is in a mixed state. However, the envelope formed by the minima of $S$ defines an upper bound for the entanglement measurement. To the best of our knowledge, this remarkable result has gone unnoticed in the literature on the matter.
@@ 2 Two-qubit states
where $z^{*}$ stands for the complex-conjugate of $z\in\mathbb{C}$. The populations (diagonal elements) and coherences (off-diagonal elements) are such that $\rho$ is self-adjoint ($\rho=\rho^{\dagger}$), normalized ($\operatorname{Tr}\rho=1$) and positive semidefinite ($\rho\geq 0$). The condition $\operatorname{Tr}\rho^{2}=1$ is necessary and sufficient for a density operator $\rho$ to represent a pure state. In turn, $\rho$ refers to mixed states if $\operatorname{Tr}\rho^{2}<1$.
@@
However, if $\Delta=0$ then $\vec{\lambda}_{1}=\vec{\lambda}_{2}=\vec{\lambda}=(\lambda,1-\lambda)$ and $\rho_{1,\operatorname{diag}}=\rho_{2,\operatorname{diag}}$. Therefore, the von Neumann entropy gives the same result for both reduced matrices $S(\rho_{1})=S(\rho_{2})$.
@@
With respect to the degree of mixing, the matrix rank of $\rho$ (the number of its nonzero eigenvalues) is quite useful since it identifies the minimum number of pure states needed to prepare the corresponding statistical mixture. In fact, this notion is directly related to the purity of $\rho$ since a pure state has rank 1 (and purity 1). So that purity less than 1 implies rank greater than 1.
@@
2.1 X–states
To construct states that satisfy the $\Delta=0$ requirement, let us first consider that only condition (3) is satisfied. In this case the density operators (1) are X–shaped,
@@ 2.1 X–states
After calculating the eigenvalues of $\rho_{X}$, see explicit expressions in Eq. (A-1) of Appendix A, we find that they are non-negative as long as the following conditions are met
@@
That is, $\rho_{X}$ represents an admissible quantum state, pure or mixed, only if the coherence amplitudes $|r_{14}|$ and $|r_{23}|$ are upper bounded by the population of the corresponding states.
@@ 2.1.1 Subsystems with equal entropy
Assuming that not only (3) but the first identity of Eq. (4) is also satisfied, states (5) will have the same population for states $|e_{1}\rangle=|00\rangle$ and $|e_{4}\rangle=|11\rangle$. As a consequence, the reduced states will have the same probability vector $\vec{\lambda}_{L}=(\lambda_{L},1-\lambda_{L})$, with $\lambda_{L}=(1-|r_{22}-r_{33}|)/2$ and $12r_{11}+r_{22}+r_{33}=$ 1. On the other hand, satisfying (3) and the second identity of Eq. (4) leads to states (5) characterized by having the same population for states $|e_{2}\rangle=|01\rangle$ and $|e_{3}\rangle=|10\rangle$. The probability vector is now $\vec{\lambda}_{R}=(\lambda_{R},1-\lambda_{R})$, with $\lambda_{R}=(1-|r_{11}-r_{44}|)/2$ and $r_{11}+2r_{22}+r_{44}=1$. Since populations satisfy $0\leq r_{kk}\leq 1$, we obtain $0\leq\lambda_{L,R}\leq\frac{1}{2}$. That is, in general, neither of the two previous cases maximizes entropy.
Reduced matrices with maximal entropy are achieved if, in addition, we make $r_{22}=r_{33}$ in $\lambda_{L}$ and $r_{11}=r_{44}$ in $\lambda_{R}$. The latter corresponds to also satisfying the complementary identity of Eq. (4) in each case. That is, only if (3) and both identities of Eq. (4) are satisfied, one gets $\lambda_{L}=\lambda_{R}=1/2$.
In Appendix A we show that once Proposition Q is satisfied, applying the above conditions to state $\rho_{X}$ leads to the Bell–states
@@
$|\beta_{3}\rangle=\frac{|e_{2}\rangle+|e_{3}\rangle}{\sqrt{2}},\qquad|\beta_{4}\rangle=\frac{|e_{3}\rangle-|e_{3}\rangle}{\sqrt{2}}.$ (8)
To give an immediate, general example, consider the convex combination
@@
This density operator is X–shaped and can accommodate two pairs of equally weighted populations. Namely, $r_{11}=r_{44}=\frac{b_{1}+b_{2}}{2}$ and $r_{22}=r_{33}=\frac{b_{3}+b_{4}}{2}$, with $r_{11}+r_{22}=\frac{1}{2}$. The other non-zero matrix elements are $r_{14}=r_{14}^{*}=\frac{b_{1}-b_{2}}{2}$ and $r_{23}=r_{23}^{*}=\frac{b_{3}-b_{4}}{2}$.
The one-qubit states that make up $\rho_{B}$ are maximally mixed $S(\rho_{B,1})=S(\rho_{B,2})=1$, regardless of $r_{11}$ and $r_{22}$. This includes the extreme cases when $\rho_{B}$ is either a Bell–state ($b_{k}=1$ for a given $k$) or an equally weighted statistical mixture of Bell–states ($b_{k}=\frac{1}{4}$ for any $k$).
An interesting particular case of the previous model is achieved by setting $b_{k}=(1+3q)/4$, and $b_{j}=(1-q)/4$ for $k=\operatorname{fixed}$, and $j\neq k$. This leads to the Werner–states [16]:
@@
These states can be generalized to the configuration
@@
where $s\leq 1+q_{\min}$ and $q_{\min}=\min\{q_{1},q_{2},q_{3},q_{4}\}$. Note that $\rho_{S}$ is another variation of $\rho_{B}$.

### v2

@@ Abstract
We investigate entanglement in two-qubit systems using a geometric representation based on the minimum of essential parameters. The latter is achieved by requiring subsystems with the same entropy, regardless of whether the state of the entire system is pure or mixed. The geometric framework is provided by a convex set $\mathcal{S}$ that forms a right-triangle, whose points are linked to just two of the coherences of the system under study. As a result, we find that optimized states of two qubits are X–shaped and host pairs of identical populations while reducing the number of coherences involved. A geometric $L$-measure of entanglement is introduced as the distance between the points in $\mathcal{S}$ that represent entangled states and the closest point that defines separable states. It is shown that $L$ reproduces the results of the Hill-Wootters concurrence $C$, so that $C$ can be interpreted as a distance-like entanglement measure. However, unlike $C$, the measure $L$ also distinguishes the rank of states. The universality of the two-qubit X–states ensures the utility of our geometric model for studying entanglement of two-qubit states in any configuration. To show the applicability of our approach far beyond time-independent cases, we construct a time-dependent two-qubit state, traced out over the complementary components of a pure tetra-partite system, and find that its one-qubit states share the same entropy. The entanglement measure results bounded from above by the envelope of the minima of such entropy.
@@ 1 Main
The entanglement in a system can be characterized and quantified from different points of view [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. However, the separability criterion introduced by Peres [19] and the Horodecki family [20] provides a powerful tool for its detection. In this work, the Peres-Horodecki (also known as positive partial transpose –PPT–) criterion is applied to construct a convex representation of two-qubit states that permits the study of entanglement in purely geometric form. It is well known that this criterion is necessary and sufficient to detect entanglement in $2\otimes 2$ and $2\otimes 3$ quantum systems [20], so we will use it to identify entanglement in two-qubit mixed states. There are several incentives to consider these states, mainly because generating pure entangled states is not an easy task, so entangled states could be prepared as statistical mixtures for purely experimental reasons, but in general they can be the result of a partial trace of the quantum state of a larger system.
@@
In general, mixed states characterize the interaction between the system and its environment, and the study of their entanglement properties is not only more complicated than that of pure states, but lesser understood. Although the first entangled system found in physics was in a pure state [2], it is in one of the Bell states [4], shortly after mixed states were identified that also disrupt the foundations of quantum theory [21]. These states, named after Werner, can be modeled using hidden variables without violating the Bell inequalities, and find applications in a diversity of quantum entanglement subjects [22, 23, 24, 25, 26, 28, 27, 29, 30, 31]. Our results include a wide class of Werner states as a particular case.
@@
We investigate the entanglement properties of two-qubit systems by using the minimum of essential parameters. The idea is to construct density operators whose reduced one-qubit states share the same entropy, regardless of whether the state of the entire system is pure or mixed. In this way, we identify X–shaped states [32, 33, 34, 35, 36, 37, 38], denoted $\rho_{X}$, that have the capacity to host pairs of identical populations while reducing the number of coherences involved. The X–states played a significant role to show that nonlocal disentanglement times are shorter than local decoherence times for spontaneous emission of two initially entangled qubits [33] (see also [34]). One of the most relevant things to note is that concurrence takes a very simple form for X–states [35, 36, 37], which justifies the extensive application of these states to study the dynamics of entanglement between two qubits [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43]. Demanding positive semidefinite, self-adjoint and normalized density operators $\rho_{X}$, and imposing the Peres-Horodecki criterion, we arrive at the system of inequalities that related matrix elements must satisfy to determine both, the eligibility of $\rho_{X}$ as a quantum state, and the possibility of finding entanglement [37]. We show that the solution of such a system defines a convex set $\mathcal{S}$ that can be divided in subsets whose points are linked to separable states and regions where entanglement may be found.
@@
Based on the convex representation provided by $\mathcal{S}$, we introduce a geometric $L$-measure of entanglement, a distance-like entanglement measure [44], that reproduces the results of the Hill-Wootters generalized concurrence $C$ [45, 46] when applied to states $\rho_{X}$. Thus, our results give $C$ a geometric meaning that is not obvious without the convex structure $\mathcal{S}$. However, unlike $C$, the measure $L$ distinguishes the rank of states with maximum entanglement. The latter is relevant to determine not only the amount of quantum correlations but rather the quantumness of the state [47], so it shows how much useful a state is for a specific quantum processing task [48, 49]. Another advantage of the geometric $L$-measure is that, with respect to other entanglement measures, it allows comparatively easy calculation.
The universality of two-qubit X–states was first explored numerically [50]. There, it is presented ‘strong numerical evidence showing that any general state can indeed be transformed to an X state of the same entanglement’. It has later been shown that X–states are in fact universal [38]. That is, given a two-qubit density operator $\rho$, a unitary transformation can be found that produces an X–shaped state $\rho_{X}$ with equal (or higher) entanglement than $\rho$. The spectrum and purity of $\rho$ are inherited unchanged to $\rho_{X}$. In this way, the geometric formulation we develop throughout this work for two-qubit X–states is useful for any configuration of the two-qubit density operators.
@@
We also study what happens to entanglement when the system depends explicitly on time. Modeling an entangled diatomic system whose one-qubit parts are placed in independent, isolated and identical electromagnetic cavities, each of which contains $n$ photons, we construct a tetra-partite pure state that continues to evolve unitarily as a pure state. However, the reduced diatomic state does not evolve in a reversible way (systems like this may show signatures of chaos [51]), which makes it an excellent subject of study to understand the behavior of entanglement in systems described by time-dependent statistical mixtures. Regardless of the dynamical law obeyed by the time-dependent diatomic state, the corresponding one-qubit states share the same von Neumann entropy $S$, which oscillates as time passes. Our results verify that finding reduced states of maximum mixing does not provide information about the entanglement (if any) of the entire system when it is in a mixed state. However, the envelope formed by the minima of $S$ defines an upper bound for the entanglement measure. To the best of our knowledge, this remarkable result has gone unnoticed in the literature on the matter.
@@ 2 Two-qubit states
where $z^{*}$ stands for the complex-conjugate of $z\in\mathbb{C}$. The populations (diagonal elements) and coherences (off-diagonal elements) are such that $\rho$ is self-adjoint ($\rho=\rho^{\dagger}$), normalized ($\operatorname{Tr}\rho=1$) and positive semidefinite ($\rho\geq 0$). The condition $\operatorname{Tr}\rho^{2}=1$ is necessary and sufficient for a density operator $\rho$ to represent a pure state. In turn, $\rho$ refers to mixed states if $\operatorname{Tr}\rho^{2}<1$. Thus, in general, matrix (1) is defined by 15 real parameters.
@@
However, if $\Delta=0$ then $\vec{\lambda}_{1}=\vec{\lambda}_{2}=\vec{\lambda}=(\lambda,1-\lambda)$ and $\rho_{1,\operatorname{diag}}=\rho_{2,\operatorname{diag}}$. In such a case the von Neumann entropy gives the same result for both reduced matrices $S(\rho_{1})=S(\rho_{2})$.
@@
With respect to the degree of mixing, the matrix rank of $\rho$ (the number of its nonzero eigenvalues) is quite useful since it identifies the minimum number of pure states needed to prepare the corresponding statistical mixture. In fact, this notion is directly related to the purity of $\rho$, $\tfrac{1}{4}\leq\eta=\operatorname{Tr}\rho^{2}\leq 1$, since a pure state has rank 1 (and purity 1). So that purity less than 1 implies rank greater than 1. In this way, the maximally mixed state $\rho_{\star}=\tfrac{1}{4}\mathbb{I}$ has rank 4 (and purity 1/4).
@@
Note that (3) removes 8 parameters from $\rho$ while (4) cancels at least one more parameter. In this way, as long as $\Delta=0$ is true, a maximum of 6 real parameters will suffice to define the quantum states $\rho$ that we study below.
2.1 Optimized states of two qubits are X–shaped
To construct states that satisfy the $\Delta=0$ requirement, let us first consider that only condition (3) is satisfied. In this case the density operators (1) require only 7 real parameters and are X–shaped,
@@ 2.1 Optimized states of two qubits are X–shaped
After calculating the eigenvalues of $\rho_{X}$, see explicit expressions in Eq. (A-1) of Appendix A, we find that they are non-negative as long as the following conditions are met [37]:
@@
That is, $\rho_{X}$ represents an admissible quantum state, pure or mixed, only if the coherence amplitudes $|r_{14}|$ and $|r_{23}|$ are upper bounded by the population of the corresponding states. Note that the phases of $r_{14}$ and $r_{23}$ play no role in characterizing $\rho_{X}$ as a positive semidefinite matrix. Nevertheless, they characterize quantum interference with respect to the corresponding distinguishable alternatives.
@@ 2.1.1 Subsystems with equal entropy
Assuming that not only (3) but the first identity of Eq. (4) is also satisfied, states (5) will have the same population for states $|e_{1}\rangle=|00\rangle$ and $|e_{4}\rangle=|11\rangle$. As a consequence, the reduced states will have the same probability vector $\vec{\lambda}_{L}=(\lambda_{L},1-\lambda_{L})$, with $\lambda_{L}=(1-|r_{22}-r_{33}|)/2$ and $12r_{11}+r_{22}+r_{33}=$ 1.
On the other hand, satisfying (3) and the second identity of Eq. (4) leads to states (5) characterized by having the same population for states $|e_{2}\rangle=|01\rangle$ and $|e_{3}\rangle=|10\rangle$. The probability vector is now $\vec{\lambda}_{R}=(\lambda_{R},1-\lambda_{R})$, with $\lambda_{R}=(1-|r_{11}-r_{44}|)/2$ and $r_{11}+2r_{22}+r_{44}=1$.
Since populations satisfy $0\leq r_{kk}\leq 1$, we obtain $0\leq\lambda_{L,R}\leq\frac{1}{2}$. That is, in general, neither of the two previous cases maximizes entropy.
Reduced matrices with maximal entropy are achieved if, in addition, we make $r_{22}=r_{33}$ in $\lambda_{L}$ and $r_{11}=r_{44}$ in $\lambda_{R}$. The latter corresponds to also satisfying the complementary identity of Eq. (4) in each case. That is, only if (3) and both identities of (4) are satisfied, one gets $\lambda_{L}=\lambda_{R}=1/2$.
The above reasoning applies to $\rho_{X}$ matrices regardless of whether they represent pure or mixed states. Therefore, when (3) and both identities in (4) hold, the density operators (5) are optimized and host pairs of identical populations. This further reduces the number of parameters needed to define $\rho_{X}$ from 7 to just 5.
With respect to the pure states represented by the optimized $\rho_{X}$ matrices, in Appendix A we show that once Proposition Q is satisfied, when (3) and both identities of (4) hold, we recover the Bell–states
@@
$|\beta_{3}\rangle=\frac{|e_{2}\rangle+|e_{3}\rangle}{\sqrt{2}},\qquad|\beta_{4}\rangle=\frac{|e_{2}\rangle-|e_{3}\rangle}{\sqrt{2}}.$ (8)
So the Bell basis elements (7)-(8) are optimized states of two qubits.
To give a more general example, consider the convex combinations
@@
This family of density operators is not only X–shaped; their members accommodate two pairs of equally weighted populations and reduce the number of coherences. Namely, the density operators $\rho_{B}$ are optimized:
$\rho_{B}=\frac{1}{2}\left(\begin{array}[]{cccc}b_{1}+b_{2}&0&0&b_{1}-b_{2}\\ 0&b_{3}+b_{4}&b_{3}-b_{4}&0\\ 0&b_{3}-b_{4}&b_{3}+b_{4}&0\\ b_{1}-b_{2}&0&0&b_{1}+b_{2}\end{array}\right).$ (10)
The one-qubit states that make up any state $\rho_{B}$ are maximally mixed $S(\rho_{B,1})=S(\rho_{B,2})=1$. This set of optimized two-qubit states includes two special configurations, the first reduces $\rho_{B}$ to a single Bell–state ($b_{k}=1$ for a given $k$), $\rho_{B}=|\beta_{k}\rangle\langle\beta_{k}|$, and the second yields an equally weighted statistical mixture of Bell–states ($b_{k}=\frac{1}{4}$ for any $k$), $\rho_{B}=\rho_{\star}$.
For states $\rho_{B}$, the coherences $r_{14}=r_{14}^{*}=b_{1}-b_{2}$ and $r_{23}=r_{23}^{*}=b_{3}-b_{4}$ are activated whenever $b_{1}\neq b_{2}$ and $b_{3}\neq b_{4}$, respectively. This property allows us to identify other interesting cases of optimized two-qubit states. For example, interference between $|e_{2}\rangle=|01\rangle$ and $|e_{3}\rangle=|10\rangle$ is not allowed if $r_{23}=0$, which can be achieved by setting $b_{3}=b_{4}=\kappa$ in (9) to get
$\left.\rho_{B}\right|_{b_{3}=b_{4}=\kappa}=b_{1}|\beta_{1}\rangle\langle\beta_{1}|+b_{2}|\beta_{2}\rangle\langle\beta_{2}|+\kappa\left(|01\rangle\langle 01|+|10\rangle\langle 10|\right),\quad b_{1}+b_{2}+2\kappa=1.$ (11)
To avoid reaching a statistical mixture of the computational basis, we will require $b_{1}\neq b_{2}$. The matrix representation is easily obtained from (10) and reads as follows
$\left.\rho_{B}\right|_{b_{3}=b_{4}=\kappa}=\left(\begin{array}[]{cccc}\frac{b_{1}+b_{2}}{2}&0&0&\frac{b_{1}-b_{2}}{2}\\ 0&\kappa&0&0\\ 0&0&\kappa&0\\ \frac{b_{1}-b_{2}}{2}&0&0&\frac{b_{1}+b_{2}}{2}\end{array}\right),\quad b_{1}\neq b_{2},\quad b_{1}+b_{2}+2\kappa=1.$ (12)
As we can see, what distinguishes this state from a statistical mixture of the basis elements is the coherence value $r_{14}=(b_{1}-b_{2})/2$. The greater the interference between $|e_{1}\rangle=|00\rangle$ and $|e_{4}\rangle=|11\rangle$, the greater the difference.
An important case of the previous model is achieved by setting $b_{k}=(1+3q)/4$, and $b_{j}=(1-q)/4$ for $k=\operatorname{fixed}$, and $j\neq k$. This leads to the Werner–states [21]:
@@
Explicitly,
$\rho_{W_{j}}=\left(\begin{array}[]{cccc}\frac{1+q}{4}&0&0&\frac{(-)^{j+1}q}{2}\\ 0&\frac{1-q}{4}&0&0\\ 0&0&\frac{1-q}{4}&0\\ \frac{(-)^{j+1}q}{2}&0&0&\frac{1+q}{4}\end{array}\right),\quad\rho_{W_{\ell}}=\left(\begin{array}[]{cccc}\frac{1-q}{4}&0&0&0\\ 0&\frac{1+q}{4}&\frac{(-)^{\ell-1}q}{2}&0\\[3.44444pt] 0&\frac{(-)^{\ell-1}q}{2}&\frac{1+q}{4}&0\\[3.44444pt] 0&0&0&\frac{1-q}{4}\end{array}\right),$ (22)
where $j=1,2$ and $\ell=3,4$. That is, the Werner–states (22) are not only X–shaped but they are also optimized in the sense discussed above.
Comparing (22) with (12), we see that $\left.\rho_{B}\right|_{b_{3}=b_{4}=\kappa}$ has the structure of the Werner–states $\rho_{W_{j}}$.
The Werner–states $\rho_{W_{k}}$ can be generalized to the configuration
@@
where $s\leq 1+q_{\min}$ and $q_{\min}=\min\{q_{1},q_{2},q_{3},q_{4}\}$. The explicit form of these optimized density operators is as follows
$\rho_{S}=\left(\begin{array}[]{cccc}\varrho_{11}&0&0&\frac{q_{1}-q_{2}}{2}\\ 0&\varrho_{22}&\frac{q_{3}-q_{4}}{2}&0\\ 0&\frac{q_{3}-q_{4}}{2}&\varrho_{22}&0\\ \frac{q_{1}-q_{2}}{2}&0&0&\varrho_{11}\end{array}\right),$ (24)
where
$\varrho_{11}=\tfrac{1}{4}\left(1+q_{1}+q_{2}-q_{3}-q_{4}\right),\qquad\varrho_{22}=\tfrac{1}{4}\left(1-q_{1}-q_{2}+q_{3}+q_{4}\right).$